OpenStudy (anonymous):

The formula for the area of a trapezoid is A= 1/2h(b1+b2). Find the value of h in terms of A, b1, and b2. Justify.

OpenStudy (anonymous):

A = (1/2) h*(b1 + b2) 2A = h*(b1 + b2) h = 2A/(b1 +b2)